OpenStudy (anonymous):

find the value for k such that the slope of the line whose equation is given will equal 8. (k-4)x+2y=10

OpenStudy (anonymous):

slope of \(ax+by=c\) is \(-\frac{a}{b}\) so set \[-\frac{k-4}{2}=8\] and solve for \(k\)

OpenStudy (anonymous):

okay so i multiplied 2 to 8 and got -k-4=16 then i added 4 and got -k=20 but i forgot how to get rid of the negative

OpenStudy (anonymous):

some mistake here

OpenStudy (anonymous):

but let me answer your question first, because you should not be confused by \[-k=20\] if \(-k=20\) then \(k=-20\) just change the sign but that is not the right answer

OpenStudy (anonymous):

oh okay. but is that how i would solve for the k ?

OpenStudy (anonymous):

the mistake you made was here \[-\frac{k-4}{2}=8\\-k-4=16\]

OpenStudy (anonymous):

the numerator carries its own invisible parentheses that is because of the order of operations you would compute in the numerator and then divide you have a choice you could write \[-\frac{k-4}{2}=8\] then multiply by 2 and get \[-(k-4)=16\\-k+4=16\\-k=12\\k=-12\]

OpenStudy (anonymous):

or you could start with \[-\frac{k-4}{2}=8\] and change it to \[\frac{k-4}{2}=-8\]

OpenStudy (anonymous):

or you could even multiply both sides by \(-2\) and get \[-\frac{k-4}{2}=8\\k-4=-16\]

OpenStudy (anonymous):

you can check that \(-12\) is the right answer by substitution

OpenStudy (anonymous):

but solving \(-x=c\) is not hard if \(-x=c\) then \(x=-c\) for example if \(-x=5\) then \(x=-5\) and if \(-x=-4\) then \(x=4\)

OpenStudy (anonymous):

so i could start the equation by plugging -12 in for k. so: (-12-4)x+2y=10 then i could slove from there

OpenStudy (anonymous):

sure you would get \(-16x+2y=10\) and if you solve for \(y\) you get \[2y=16x+10\] or \[y=8x+5\] a line with slope \(8\) as required

OpenStudy (anonymous):

okay thank you!(: